solve the quadatric equations by factoring
(x-1)^2=7

Answer by stanbon(75887)

(x-1)^2=7
--------------
(x-1)^2-7 = 0
(x-1-sqrt(7))(x-1+sqrt(7))
------
x = 1+sqrt7 or x = 1-sqrt(7)
=================
Cheers,
Stan H.